OpenStudy (anonymous):

Solve this system of equations using the elimination method: y = 2x + 8 4x + 2y = 10

hartnn (hartnn):

subtract 2x from both sides of first equation

OpenStudy (anonymous):

y - 2x = 8?

OpenStudy (anonymous):

Sorry to bother you again btw.

hartnn (hartnn):

yes and no problem :) now multiply 2 on both side

OpenStudy (anonymous):

Both sides of both equations?

hartnn (hartnn):

-2x +y =8 here, multiply both sides by 2 thats because , to eliminate, we try to make the co-efficients of the variable equal

OpenStudy (anonymous):

mkay. -4x + 2y = 16

hartnn (hartnn):

now add the 2 equations! -4x+2y =16 4x+2y = 10 and observe that 'x' gets eliminated :)

OpenStudy (anonymous):

Oh, okay. So 2y = 26 2y/2 = 26/2 y = 13

hartnn (hartnn):

2y+2y = 4y :P

OpenStudy (anonymous):

Oh. :/ 4y = 26 4y/4 = 26/4 y = 6.5

hartnn (hartnn):

correct :) now find x

OpenStudy (anonymous):

So now.... y = 2x + 8 6.5 = 2x + 8 -8 = -8 -1.5 = 2x -1.5/2 = 2x/2 -0.75 = x

OpenStudy (anonymous):

Thank you again. :-D

hartnn (hartnn):

you're good at this! welcome ^_^